③ The transfer function T(s): represents a GPI> System with a unit slep response of 25%overshoot a setling iime T isec. Th ess -o. Determine the constants K ,k2-_-)k3= nd e steady sae error
Answer #1

Ans)

Given T(s)=rac{K_{1}}{s^{2}+K_{2}s+K_{3}}

The steady state error is given as

ess-lim sE(s) = lim sR(s)(1-T(s))

For step input R(s)=1/s

e_{ss}=lim_{s ightarrow 0}sE(s)=lim_{s ightarrow 0}srac{1}{s}(1-T(s))

e_{ss}=lim_{s ightarrow 0}(1-T(s))=lim_{s ightarrow 0}(1-rac{K_{1}}{s^{2}+K_{2}s+K_{3}})

K1 ess (1 0.1

rac{K_{1}}{K_{3}}=1-0.1=0.9

rac{K_{1}}{K_{3}}=0.9

From the denominator of transfer function i.e s^{2}+K_{2}s+K_{3} comparing with standard second order equation i.e

s^{2}+2zeta w_{n}s+w^{2}_{n}

w^{2}_{n}=K_{3}Rightarrow w_{n}=sqrt{K_{3}}

2zeta w_{n}=K_{2}

zeta =rac{K_{2}}{2w_{n}}=rac{K_{2}}{2sqrt{K_{3}}}

Given settling time

T_{s}=rac{4}{zeta w_{n}}=1

zeta w_{n}=4

rac{K_{2}}{2sqrt{K_{3}}}*sqrt{K_{3}}=4

rac{K_{2}}{2}=4

mathbf{K_{2}=8}

Given overshoot

o/oM_{p}=25: : o/o

M_{p}=0.25=e^{-rac{zeta pi }{sqrt{1-zeta ^{2}}}}

-rac{zeta pi }{sqrt{1-zeta ^{2}}}=ln(0.25)=-1.386

rac{zeta pi }{sqrt{1-zeta ^{2}}}=1.386

zeta^{2} pi^{2}=1.386^{2}(1-zeta ^{2})

zeta^{2}(1.386^{2}+ pi^{2})=1.386^{2}

zeta=sqrt{rac{1.386^{2}}{(1.386^{2}+ pi^{2})}}=0.404

zeta=0.404=rac{K_{2}}{2sqrt{K_{3}}}

using K_{2}=8

zeta=0.404=rac{8}{2sqrt{K_{3}}}

K_{3}=(rac{8}{2*0.404})^{2}=98.17

mathbf{K_{3}=98.17}

Now using rac{K_{1}}{K_{3}}=0.9

K_{1}=0.9K_{3}=0.9*98.17=88.35

mathbf{K_{1}=88.35}mathbf{K_{2}=8}mathbf{K_{3}=98.17}